How many pounds of cashew nuts selling for 90 cents a pound must be mixed with 30 pounds of peanuts selling at 40 cents to make a mixture which sell for 65 cents a pound?
Thanks!

Answer by checkley79(3341)

.90X+.40*30=.85(30+X)
.90X+12=25.50+.85X
.90X-.85X=25.50-12
.05X=13.50
X=13.50/.05
X=270 POUNDS OF CAHSEWS ARE USED.
PROOF:
.90*270+.40*30=.85(30+270)
243+12=.85*300
255=255